# On the concentration distribution in turbulent thermals

### v1

@@ Abstract
Turbulent thermals emerge in a wide variety of geophysical and industrial flows, such as atmospheric cumulus convection and pollutant dispersal in oceans and lakes. They form when a volume of buoyant fluid, arising from thermal or compositional density differences, is instantly released from a source. As they rise or sink, heat and mass transfers are likely to occur with the surrounding environment, spanning multiple scales from macroscopic entrainment of ambient fluid to microscopic diffusive processes. Although turbulent thermals are typically investigated through their integral properties, mixing processes specifically depend on the internal distribution of concentration or temperature. Here, we use laboratory fluid dynamics experiments and direct numerical simulations to investigate the mixing of a passive scalar in turbulent thermals with large Reynolds numbers. We track the evolution of the concentration field, computing its moments and the probability density function. The concentration distribution exhibits self-similarity over time, except at high concentrations, possibly because of the presence of undiluted cores. These distributions are well approximated by an exponential probability density function. In the investigated range, we find that concentration distributions are largely independent of the Péclet and Reynolds numbers.
keywords:
1 Introduction
A finite volume of buoyant fluid evolving under the influence of gravity in an otherwise quiescent environment forms what is called a turbulent thermal when the flow is turbulent. The term thermal originates from the terminology used by glider pilots, referring to a volume of warm air rising through the atmosphere (Scorer, 1957; Woodward, 1958). Turbulent thermals are fundamental elements of atmospheric convection and have been extensively studied by both the fluid dynamics and atmospheric science communities (Batchelor, 1954; Blyth et al., 2005; Yano, 2014). Yet the interest in turbulent thermals also extends to other contexts. Flows sharing strong similarities with atmospheric thermals can also develop in liquid environments, without being necessarily driven by temperature-induced density variations. The nature of the buoyancy forces — whether due to temperature variations, compositional differences, or the presence of multiple phases with differing densities — appears to be of secondary importance for the overall behavior of the thermal (Batchelor, 1954; Turner, 1979; Bush et al., 2003; Landeau et al., 2014).
Field observations and laboratory experiments show that turbulent thermals expand and decelerate as they rise, as a result of the gradual entrainment of ambient fluid (Batchelor, 1954; Scorer, 1957; Woodward, 1959). This is consistent with dimensional analysis (Batchelor, 1954), which predicts that far from the source the thermal radius $r$, velocity $w$, and reduced gravitational acceleration (or local buoyancy) $g^{\prime}$ evolve with distance $z$ and time $t$ as
@@ 1 Introduction
It is instructive to analyze the implications of equations 1-3 for the evolution with $z$ of the different terms of the momentum conservation equation. Assuming that velocity fluctuations vary over a length scale $\sim r$, a scale analysis of the different terms of the momentum conservation equation shows that inertia scales as $\mathbf{v}\cdot\mathbf{\nabla}\mathbf{v}\sim\mathcal{B}_{0}z^{-3}$, buoyancy $g^{\prime}\sim\mathcal{B}_{0}z^{-3}$, and viscous forces $\nu\mathbf{\nabla}^{2}\mathbf{v}\sim\nu\mathcal{B}_{0}^{1/2}z^{-3}$ ($\nu$ denotes kinematic viscosity). The ratio of inertia to viscous forces, which defines an instantaneous Reynolds number $Re$, is constant in time (Scorer, 1957; Johari, 1992; Morton et al., 1994) and given by
@@
The fact that the instantaneous Reynolds and Péclet numbers are constant suggests that turbulent thermals can evolve in a self-similar way when far from the source, in the sense that velocity and concentration or temperature would each have similar distributions at different times or distances from the source. This is consistent with early experimental determinations of the mean velocity field (Woodward, 1959), and with experimental observations showing that the shape of thermals is often self-preserved (Scorer, 1957).
@@
However, some experimental observations argue against true self-similarity. Bond & Johari (2010) found that the vorticity within turbulent thermals tends to be localized in thin vortex cores, the width of which increases more slowly than the thermal radius. This was confirmed by Zhao et al. (2013), who also found that the maximum concentration within the thermal decreases at a much slower rate than predicted for the average concentration according to 3. Note also that the scaling laws 1-3, which imply constant $Re$ and $Pe$, assume that the total buoyancy of the thermal is conserved. Experiments and numerical simulations show that detrainment (i.e loss of mass and buoyancy) is small (Johari, 1992; Lecoanet & Jeevanjee, 2019; Morrison, 2025) but not inexistent (an order of magnitude smaller than entrainment according to Lecoanet & Jeevanjee (2019)). This implies that the instantaneous $Re$ and $Pe$ decrease with time, albeit slightly, which does not allow perfect self-similarity.
@@
The question of self-similarity is also related to the dynamics of mixing and homogenization of thermal or compositional heterogeneities, which sets the distribution of temperature or concentration within the thermal. The process of mixing involves the irreversible transformation of initially segregated phases into a homogeneous state (e.g. Villermaux, 2019). Unlike confined mixtures, a state of complete homogeneity can never be achieved in an expanding turbulent thermal due to the continuous entrainment of the surrounding fluid. The competition between homogenization, driven by stirring motions, and the production of compositional heterogeneities through the entrainment of surrounding fluid determines the degree of mixing within a thermal. A self-similar evolution would require that these two processes are in balance.
@@
To our knowledge, the statistical distribution of the concentration or temperature field in a turbulent thermal has not yet been investigated, at least not in any systematic way. Numerical simulations of well-developed turbulent thermals (Lai et al., 2015; Lecoanet & Jeevanjee, 2019; McKim et al., 2020; Orlandi & Carnevale, 2020; Vybhav & Ravichandran, 2022; Morrison et al., 2022; Middleton & Mumford, 2023; Morrison et al., 2023; Morrison, 2025) have predominantly investigated their global behavior in terms of the effects of moisture content on entrainment rates or the influence of stratification on thermal rise. Previous experimental work has also examined concentration measurements in puffs (Yee et al., 1994; Ghaem-Maghami & Johari, 2006), plumes (Yee et al., 1995), and thermals (Bat et al., 1984; Marcus & Bell, 1992; Zhao et al., 2013). However, they mostly provide global concentration measurements (Marcus & Bell, 1992; Bond & Johari, 2005, 2010; Zhao et al., 2013).
Our main goal here is to quantify the time-evolution of the distribution of the concentration or temperature field, as measured by their Probability Density Function (PDF), using laboratory experiments and numerical simulations at different Schmidt numbers $Sc=\nu/D$. This provides a way of quantifying the approach to self-similarity as well as possible departures from self-similarity. It is also a first step towards a more detailed analysis of the dynamics of mixing within a turbulent thermal. The turbulent thermal, with its well-defined flow structure and its self-similar characteristics, is an interesting prototype for studying dispersion and mixing in an unconfined environment.
@@
In this paper, we first revisit the dimensional analysis of turbulent thermals and propose predictions for the time evolution of the moments of the concentration field. In Section 3, we describe our experimental setup and numerical simulations, which were performed at Reynolds numbers up to $Re=10^{4}$. Our direct 3D numerical simulations match the experimental Reynolds number but use a significantly lower Schmidt number ($Sc=1$). We analyze the influence of $Re$ and $Sc$ on the evolution and distribution of the concentration field. Section 4 provides a qualitative description of turbulent thermals in experiments and simulations, including the definition of thermal boundaries and their dependence on concentration and vorticity thresholds. Section 5 presents a quantitative analysis of concentration distributions and moments, which suggests that the concentration field follows an exponential distribution far from the source. Finally, we discuss the effects of diffusivity ($Sc$) on concentration evolution in Section 5.4.
2 Dimensional analysis
2.1 Bulk properties
@@
Previous experiments have shown that a turbulent thermal continuously entrains ambient fluid as it falls or rises (Batchelor, 1954; Scorer, 1957; Turner, 1979; Zhao et al., 2013). As a result, the volume $V$ of the thermal increases over time, while both its average concentration $\langle c\rangle$ and mean density difference $\Delta\rho$ relative to the surrounding fluid decrease (e.g. Batchelor, 1954; Morton et al., 1956). Since the loss of material from the thermal is small (Lecoanet & Jeevanjee, 2019), we assume that the mass of solute within the thermal, defined as $\mathcal{M}=\langle c\rangle V$, is conserved and equal to its initial value
@@
In the Boussinesq approximation limit, a thermal is fully defined by $V_{0}$, $\mathcal{B}_{0}$, $\mathcal{M}_{0}$, and by the transport properties $\nu$ and $D$. As shown by Batchelor (1954), dimensional analysis provides valuable predictions for the bulk properties of a thermal. For instance, the vertical position $z_{c}$ of the center of mass of the thermal is a function of $t$, $V_{0}$, $\mathcal{B}_{0}$, $\mathcal{M}_{0}$, $\nu$, and $D$. This gives a set of seven quantities with three independent units. According to Vaschy-Buckingham’s theorem, only four independent dimensionless quantities can be formed from these quantities, one possible set being
@@
In these equations, $V_{0}$, $\mathcal{B}_{0}$, $\mathcal{M}_{0}$, and $Re$, $Pe$ are conserved, time-independent quantities that do not depend on the release conditions of the dense fluid. In contrast, $t_{g}$ is influenced by the specifics of the initial conditions. When far enough from the release point, the thermal is expected to become independent of its initial conditions (Batchelor, 1954). In this limit, $r$, $w$, and $\langle c\rangle$ become independent of $t_{g}$. In addition, laboratory experiments and field observations of atmospheric thermals (e.g. Woodward, 1958, 1959) and volcanic clouds (e.g. Terada & Ida, 2007) have shown no measurable effect of $Re$ and $Pe$ on the time evolution of thermal size and velocity. When $t\gg t_{g}$, these assumptions imply that (Batchelor, 1954; Morton et al., 1956)
@@
where $\alpha$ is the entrainment coefficient (Morton et al., 1956).
2.2 Moments of concentration
Using dimensional analysis along the same lines as in section 2.1, the moments of order $k$ $\langle c^{k}\rangle$ of the concentration field can be expressed as
@@ 2.2 Moments of concentration
As explained in section 2.1, we expect that the thermal becomes independent of its initial condition when far enough from the release point (Batchelor, 1954). In this limit, when $t\gg t_{g}$, the moments become independent of $t_{g}$, which gives
@@
This form has the inconvenience of artificially introducing parameters related to the source ($c_{0}$ and $t_{g}$) but will prove more convenient for comparison with experimental and numerical data (section 5).
The dimensional analysis on the moments of concentration (Eq. 20) implies that the shape of the PDF of $c$ may vary with $Re$ and $Pe$. As a result, we aim to evaluate the predictions for the time evolution of the moments of $c$ and to investigate the effect of $Re$ and $Pe$ on the shape of its distribution. Although it has been convincingly argued that the bulk properties of the thermal are expected to be independent of viscosity and diffusivity (Taylor, 1946; Batchelor, 1954; Scorer, 1957; Morton et al., 1956), this has yet to be tested for the concentration distribution.
3 Experiments and numerical simulations
3.1 Experiments
In our experiments, we release a volume of an aqueous solution of sodium chloride (NaCl) into a $30\times 30\times 100$ cm tank containing fresh water (figure 1). The NaCl solution is initially contained in aluminum tubes of varying diameter and height $s_{0}=15$, $30$, $45$, and $60$ mm, corresponding to released volumes $V_{0}=2.7$, $21.2$, $71.6$, and $169.6$ mL, respectively. We use the diameter $d$ of the equivalent sphere of volume $V_{0}$, i.e. $d=(3/2)^{1/3}s_{0}$, as a measure of the initial size of the released volume. We seal the tube at both ends with latex membranes stretched over the openings and proceed to fill it with the NaCl solution. We submerge the tube in the tank to a depth of approximately 15 cm below the surface, and we delicately pierce the upper membrane with a needle to release any trapped air bubbles. A negligible amount of fluid is drawn out of the tube during the operation due to the retraction of the membrane. After a minute of rest, we release a needle previously held in place by an electromagnet and abruptly puncture the lower membrane, which retracts rapidly. The fluid contained in the tube is quickly released due to the density difference with the ambient fluid. For the 15 mm tube, we pierce both membranes simultaneously, as, in this case, the retraction of the upper membrane draws a non-negligible amount of fluid out of the tube. As a result, dense fluid is released from the top of the tube, significantly altering the initial conditions of the fluid release.
@@ 3.1 Experiments
We obtain the concentration field of a fluorescent dye (Rhodamine 6G) initially mixed with the NaCl solution using Laser-Induced Fluorescence (LIF). We illuminate the tank from below with a vertical laser sheet (532 nm) produced by a continuous 400 mW Nd:YAG laser coupled with a cylindrical lens (figure 1). We image the LIF field using a PCO Edge 5.5 camera ($2560\times 2160$ pixels, 20 Hz, 16-bit dynamic range) fitted with a Nikon Nikkor 50 mm f/1.2 lens. We use a high-pass filter ($>540$ nm) to remove the laser wavelength and isolate the emission signal of Rhodamine 6G (between 500 and 700 nm). We spatially calibrate the images using a grid, remove the background, and correct for laser sheet heterogeneities. Finally, we obtain the concentration field using a calibration between light intensity and concentration (Appendix A).
@@
Our experiments are divided into four categories, (A) to (D), depending on the tube used to release the dense fluid (Table 1). For each tube, we choose the NaCl concentration such that the Reynolds number is maximized while keeping a concentration low enough to prevent significant optical distortion when the thermal enters the field of view of the camera. We measure the density of the NaCl solutions using an Anton Paar 35 Basic densitometer, which has an accuracy of approximately $0.1\%$. We estimate viscosities and refractive indices from data compiled by Haynes (2014).
[…]
In each configuration, we conducted and analyzed three experiments to estimate flow variability. The unsteady nature of the flow, as well as the coupling between buoyancy and mean flow, makes the early evolution of the thermal sensitive to the initial conditions resulting from puncturing the membrane that seals the tube. The lower membrane does not always retract in the same way, thereby affecting the initial conditions. In Section 5, we will show the evolution of concentration moments and PDF of concentration of the average of the three experiments for each Reynolds number.
# $s_{0}$ (mm) $d$ (mm) $\Delta\rho_{0}/\rho_{a}$ $\mathcal{B}_{0}$ ($\mathrm{m}^{4}\mathrm{/}\mathrm{s}\mathrm{{}^{2}}$) $\mu_{0}/\mu_{a}$ $Re$ $Pe$
[…]

### v2

@@ Abstract
Turbulent thermals emerge in a wide variety of geophysical and industrial flows, such as atmospheric cumulus convection and pollutant dispersal in oceans and lakes. When a buoyant fluid mass rises, or sinks, heat and mass transfers occur by the engulfment of the fresh surrounding fluid inside the thermal - a process that spans over multiple scales from macroscopic entrainment of ambient fluid to microscopic diffusive processes. Turbulent thermals are typically investigated through their integral properties (radius, depth, entrainment rate). However, mixing processes depend on the internal distribution of concentration or temperature inside a thermal, which remains poorly constrained. Here, we use laboratory fluid dynamics experiments and direct numerical simulations to investigate the mixing of a passive scalar in turbulent thermals with large Reynolds numbers. We track the evolution of the concentration field, computing its moments and the probability density function. The concentration distribution exhibits self-similarity over time, except at high concentrations, possibly because of the presence of undiluted cores. These distributions are well approximated by an exponential probability density function. Although diffusion has a strong effect on the spatial structure of the concentration field, we observe no significant effect of diffusivity on the concentration distributions in the investigated range of Péclet numbers.
I Introduction
A finite volume of buoyant fluid evolving under the influence of gravity in an otherwise quiescent environment forms what is called a turbulent thermal when the flow is turbulent. The term thermal originates from the terminology used by glider pilots, referring to a volume of warm air rising through the atmosphere [35, 43]. Turbulent thermals are fundamental elements of atmospheric convection and have been extensively studied by both the fluid dynamics and atmospheric science communities [2, 4, 45]. Yet the interest in turbulent thermals also extends to other contexts. Flows sharing strong similarities with atmospheric thermals can also develop in liquid environments, without being necessarily driven by temperature-induced density variations. The nature of the buoyancy forces — whether due to temperature variations, compositional differences, or the presence of multiple phases with differing densities — appears to be of secondary importance for the overall behavior of the thermal [2, 39, 9, 18].
Field observations and laboratory experiments show that turbulent thermals expand and decelerate as they rise, as a result of the gradual entrainment of ambient fluid [2, 35, 44]. This is consistent with dimensional analysis [2], which predicts that far from the source the thermal radius $r$, velocity $w$, and reduced gravitational acceleration (or local buoyancy) $g^{\prime}$ evolve with distance $z$ and time $t$ as
@@ I Introduction
It is instructive to analyze the implications of equations 1-3 for the evolution with $z$ of the different terms of the momentum conservation equation. Assuming that velocity fluctuations vary over a length scale $\sim r$, a scale analysis of the different terms of the momentum conservation equation shows that inertia scales as $\mathbf{v}\cdot\mathbf{\nabla}\mathbf{v}\sim\mathcal{B}_{0}z^{-3}$, buoyancy $g^{\prime}\sim\mathcal{B}_{0}z^{-3}$, and viscous forces $\nu\mathbf{\nabla}^{2}\mathbf{v}\sim\nu\mathcal{B}_{0}^{1/2}z^{-3}$ ($\nu$ denotes kinematic viscosity). The ratio of inertia to viscous forces, which defines an instantaneous Reynolds number $Re$, is constant in time [35, 16, 29] and given by
@@
The fact that the instantaneous Reynolds and Péclet numbers are constant suggests that turbulent thermals can evolve in a self-similar way when far from the source, in the sense that velocity and concentration or temperature would each have similar distributions at different times or distances from the source. This is consistent with early experimental determinations of the mean velocity field [44], and with experimental observations showing that the shape of thermals is often self-preserved [35].
@@
However, some experimental observations argue against self-similarity. Bond & Johari (2010) [6] found that the vorticity within turbulent thermals tends to be localized in thin vortex cores, the width of which increases more slowly than the thermal radius. This was confirmed by Zhao et al. (2013) [48], who also found that the maximum concentration within the thermal decreases at a much slower rate than predicted for the average concentration according to Eq. (3). Note that the scaling laws (Eq.(1-3)), which imply constant $Re$ and $Pe$, assume that the total buoyancy of the thermal is conserved. Experiments and numerical simulations show that detrainment (i.e, loss of mass and buoyancy) is small [16, 22, 28] but not inexistent (an order of magnitude smaller than entrainment according to [22]). This implies that the instantaneous $Re$ and $Pe$ decrease with time, albeit slightly, which does not point toward asymptotic self-similarity.
@@
The question of self-similarity of the concentration is also related to the dynamics of mixing and homogenization of thermal or compositional heterogeneities, which sets the distribution of temperature or concentration within the thermal. The process of mixing involves the irreversible transformation of initially segregated phases into a homogeneous state [e.g. 41]. Unlike confined mixtures, a state of complete homogeneity can never be achieved in an expanding turbulent thermal due to the continuous entrainment of the surrounding fluid. The competition between homogenization, driven by local stirring motions, and the production of compositional heterogeneities through the large-scale entrainment of surrounding fluid determines the degree of mixing within a thermal. It remains an open question whether these two processes are in dynamic equilibrium.
@@
To our knowledge, the statistical distribution of the concentration or temperature field in a turbulent thermal has not yet been investigated. Numerical simulations of well-developed turbulent thermals [17, 22, 24, 31, 42, 27, 25, 26, 28] have predominantly investigated their global behavior in terms of the effects of moisture content on entrainment rates or the influence of stratification on thermal rise. Previous experimental work has also examined concentration measurements in puffs [46, 13], plumes [47], and thermals [1, 23, 48]. However, they mostly provide average concentration measurements [23, 5, 6, 48]. Here, we quantify the time-evolution of the distribution of the concentration or temperature fields, as measured by their Probability Density Function (PDF), using both laboratory experiments and numerical simulations at different Schmidt numbers $Sc=\nu/D$. Our direct 3D numerical simulations match the experimental Reynolds number but use a significantly lower Schmidt number ($Sc=1$). We analyze the influence of $Re$ and $Sc$ on the evolution and distribution of the concentration field.
@@
The paper is organized as follows. We first recall the dimensional analysis of turbulent thermals. We propose predictions for the time evolution of the moments of the concentration field. In Section III, we describe the experimental setup and numerical simulations. Section IV provides a qualitative description of turbulent thermals in experiments and simulations, including the definition of thermal boundaries and their dependence on concentration and vorticity thresholds. Section V presents a quantitative analysis of concentration distributions and moments, which suggests that the concentration field follows an exponential distribution far from the source. Finally, we discuss the effects of diffusivity ($Sc$) on concentration evolution in Section V.4.
II Dimensional analysis
II.1 Bulk properties
@@
Previous experiments have shown that a turbulent thermal continuously entrains ambient fluid as it falls or rises [2, 35, 39, 48]. As a result, the volume $V$ of the thermal increases over time, while both its average concentration $\langle c\rangle$ and mean density difference $\Delta\rho$ relative to the surrounding fluid decrease [e.g. 2, 30]. Since the loss of material from the thermal is small [22], we assume that the mass of solute within the thermal, defined as $\mathcal{M}=\langle c\rangle V$, is conserved and equal to its initial value
@@
In the Boussinesq approximation limit, a thermal is fully defined by $V_{0}$, $\mathcal{B}_{0}$, $\mathcal{M}_{0}$, and by the transport properties $\nu$ and $D$. As shown by [2], dimensional analysis provides valuable predictions for the bulk properties of a thermal. For instance, the vertical position $z_{c}$ of the center of mass of the thermal is a function of $t$, $V_{0}$, $\mathcal{B}_{0}$, $\mathcal{M}_{0}$, $\nu$, and $D$. This gives a set of seven quantities with three independent units. According to Vaschy-Buckingham’s theorem, only four independent dimensionless quantities can be formed from these quantities, one possible set being
@@
In these equations, $V_{0}$, $\mathcal{B}_{0}$, $\mathcal{M}_{0}$, and $Re$, PePe are conserved, time-independent quantities that do not depend on the release conditions of the dense fluid. In contrast, $t_{g}$ is influenced by the specifics of the initial conditions. When far enough from the release point, the thermal is expected to become independent of its initial conditions [2]. In this limit, $r$, $w$, and $\langle c\rangle$ become independent of $t_{g}$. In addition, laboratory experiments and field observations of atmospheric thermals [e.g. 43, 44] and volcanic clouds [e.g. 37] have shown no measurable effect of $Re$ and $Pe$ on the time evolution of thermal size and velocity. When $t\gg t_{g}$, these assumptions imply that [2, 30]
@@
where $\alpha$ is the entrainment coefficient [30].
II.2 Moments of concentration
Using dimensional analysis along the same lines as in section II.1, the moments of order $k$ $\langle c^{k}\rangle$ of the concentration field can be expressed as
@@ II.2 Moments of concentration
As explained in section II.1, we expect that the thermal becomes independent of its initial condition when far enough from the release point [2]. In this limit, when $t\gg t_{g}$, the moments become independent of $t_{g}$, which gives
@@
This form has the inconvenience of artificially introducing parameters related to the source ($c_{0}$ and $t_{g}$) but will prove more convenient for comparison with experimental and numerical data (section V).
The dimensional analysis on the moments of concentration (Eq. 20) implies that the shape of the PDF of $c$ may vary with $Re$ and $Pe$. As a result, we aim to evaluate the predictions for the time evolution of the moments of $c$ and to investigate the effect of $Re$ and $Pe$ on the shape of its distribution. Although it has been convincingly argued that the bulk properties of the thermal are expected to be independent of viscosity and diffusivity [36, 2, 35, 30], this has yet to be tested for the concentration distribution.
III Experiments and numerical simulations
III.1 Experiments
In our experiments, we release a volume of an aqueous solution of sodium chloride (NaCl) into a $30\times 30\times 100$ cm tank containing fresh water (Fig. 1). The NaCl solution is initially contained in aluminum tubes of varying diameter and height $s_{0}=15$, $30$, $45$, and $60$ mm, corresponding to released volumes $V_{0}=2.7$, $21.2$, $71.6$, and $169.6$ mL, respectively. We use the diameter $d$ of the equivalent sphere of volume $V_{0}$, i.e. $d=(3/2)^{1/3}s_{0}$, as a measure of the initial size of the released volume. We seal the tube at both ends with latex membranes stretched over the openings and proceed to fill it with the NaCl solution. We submerge the tube in the tank to a depth of approximately 15 cm below the surface, and we delicately pierce the upper membrane with a needle to release any trapped air bubbles. A negligible amount of fluid is drawn out of the tube during the operation due to the retraction of the membrane. After a minute of rest, we release a needle previously held in place by an electromagnet and abruptly puncture the lower membrane, which retracts rapidly. The fluid contained in the tube is quickly released due to the density difference with the ambient fluid. For the 15 mm tube, we pierce both membranes simultaneously, as, in this case, the retraction of the upper membrane draws a non-negligible amount of fluid out of the tube. As a result, dense fluid is released from the top of the tube, significantly altering the initial conditions of the fluid release.
@@ III.1 Experiments
We obtain the concentration field of a fluorescent dye (Rhodamine 6G) initially mixed with the NaCl solution using Laser-Induced Fluorescence (LIF). We illuminate the tank from below with a vertical laser sheet (532 nm) produced by a continuous 400 mW Nd:YAG laser coupled with a cylindrical lens (Fig. 1). We image the LIF field using a PCO Edge 5.5 camera ($2560\times 2160$ pixels, 20 Hz, 16-bit dynamic range) fitted with a Nikon Nikkor 50 mm f/1.2 lens. We use a high-pass filter (>540>540 nm) to remove the laser wavelength and isolate the emission signal of Rhodamine 6G (between 500 and 700 nm). We spatially calibrate the images using a grid, remove the background, and correct for laser sheet heterogeneities. Finally, we obtain the concentration field using a calibration between light intensity and concentration (Appendix A). The experimental values of concentration are limited by the dynamic range of the camera (16-bit resolution), and they contain a maximum of 5 orders of magnitude in intensity variation.
@@
Our experiments are divided into four categories, (A) to (D), depending on the tube used to release the dense fluid (Table 1). For each tube, we choose the NaCl concentration such that the Reynolds number is maximized while keeping a concentration low enough to prevent significant optical distortion when the thermal enters the field of view of the camera. We measure the density of the NaCl solutions using an Anton Paar 35 Basic densitometer, which has an accuracy of approximately $0.1\%$. We estimate viscosities and refractive indices from data compiled by Haynes [14].
Table 1 gives the ratio $\Delta\rho_{0}/\rho_{a}$ selected for the four configurations corresponding to different tube diameters $s_{0}$ and equivalent sphere diameter $d$. The range of initial density differences corresponds to a range of total buoyancy $\mathcal{B}_{0}$ and translates into a viscosity range of $\mu_{0}=1.03-1.60~$\mathrm{mPa}\text{\,}\mathrm{s}$\$ and a refractive index range of 1.3361.3701.336-1.370. In comparison, the viscosity of fresh water is $\mu_{a}=1.005~$\mathrm{mPa}\text{\,}\mathrm{s}$\$ and its refractive index is $1.333$. However, it should be noted that the cloud formed by the NaCl solution is significantly diluted when it enters the field of view of the camera, resulting in variations of viscosity and refractive index that are significantly smaller than those suggested by the values mentioned above. The maximum viscosity corresponding to the maximum salt concentration of the thermal (see Fig. 5b), is up to 3% larger than the viscosity of water (1.6%, 2.3%, 3%, 3% for A, B, C, D cases). The average viscosity in the diluted thermal when it enters the field of view is then close to the water viscosity. Then, we neglect the effect of viscosity on the long-term evolution of the thermal and calculate the $Re$ using the viscosity of the ambient fluid (see Eq. 8).
Péclet numbers are calculated from equation (8), using the diffusivity of Rhodamine 6G in water, estimated at $D^{rhod}=$4.0\text{\times}{10}^{-10}$~${\mathrm{m}}^{2}\text{\,}{\mathrm{s}}^{-1}$\$ [12], for $Pe$. The diffusivity of sodium chloride, $D^{salt}=$1.5\text{\times}{10}^{-9}$~${\mathrm{m}}^{2}\text{\,}{\mathrm{s}}^{-1}$\$, is approximately 4 times larger than the diffusivity of Rhodamine 6G. As a result, we expect the salt concentration field to differ from the Rhodamine 6G concentration field measured with LIF. However, both diffusion coefficients give a large Schmidt number $Sc$ (2500 and 6700, respectively).
In each configuration, we conducted and analyzed three experiments to estimate flow variability. The unsteady nature of the flow, as well as the coupling between buoyancy and mean flow, makes the early evolution of the thermal sensitive to the initial conditions resulting from puncturing the membrane that seals the tube. The lower membrane does not always retract in the same way, thereby affecting the initial conditions. The measured concentration is normalized by the initial concentration $c_{0}=c(t=0)$ in the tube. In Section V, we will show the evolution of concentration moments and PDF of concentration of the average of the three experiments for each Reynolds number.
# $s_{0}$ ($m$) $d$ ($m$) $\Delta\rho_{0}/\rho_{a}$ $\mathcal{B}_{0}$ ($\mathrm{m}^{4}\mathrm{/}\mathrm{s}\mathrm{{}^{2}}$) $\mu_{0}/\mu_{a}$ ReRe PePe
(A) $15$ $17.1$ $0.156$ $4.1\text{\times}{10}^{-6}$ $1.62$ $2012$ $5.04\text{\times}{10}^{6}$
(B) $45$ $51.5$ $0.017$ $1.2\text{\times}{10}^{-5}$ $1.05$ $3452$ $8.64\text{\times}{10}^{6}$
(C) $30$ $34.3$ $0.084$ $1.8\text{\times}{10}^{-5}$ $1.25$ $4175$ $1.04\text{\times}{10}^{7}$
(D) $60$ $68.7$ $0.013$ $2.2\text{\times}{10}^{-5}$ $1.04$ $4646$ $1.16\text{\times}{10}^{7}$